# BCTR: Bidirectional Conditioning Transformer for Scene Graph Generation

### v2

@@ 4.3 Random Feature Alignment
Random Feature Alignment As illustrated in Fig. 1 (a), the class imbalance in SGG datasets makes it challenging for BCG to capture feature interaction patterns for tail relationships from rare samples during training. Additionally, the learned feature interactions perform poorly on unseen relationships during inference. To address these issues, we propose RFA, which aligns the SGG feature space with that of a pre-trained vision-language model. In this semantically aligned feature space, BCG can effectively learn feature interaction patterns during training, and the learned patterns generalize better to unseen relationships.
@@
Specifically, the input image is fed into the CLIP visual encoder to obtain $V_{\text{clip}}$. Since CLIP is trained with image captions, the encoder features $V_{\text{clip}}$ are inclined to capture the gist of the image. However, images often contain multiple relationships. Simply aligning the decoder’s features $Q_{p}^{\text{end}}$ and $Q_{e}^{\text{end}}$ with the CLIP feature $V_{\text{clip}}$ may compromise the feature diversity of SGG. To preserve this diversity, we randomly select a subset of the decoder’s features before each alignment, ensuring that the distilled features retain diversity while being aligned with CLIP. The overall masked distillation process is as follows:
$\mathcal{L}_{mimic}=\left|V_{clip}-\frac{1}{N_{p}}\sum_{i=1}^{N_{p}}\bar{Q}_{p}^{end}\right|+\left|V_{clip}-\frac{1}{N_{e}}\sum_{i=1}^{N_{e}}\bar{Q}_{e}^{end}\right|,$
where $\bar{Q}_{p}^{\text{end}}$ and $\bar{Q}_{e}^{\text{end}}$ are randomly sampled from $Q_{p}^{\text{end}}$ and $Q_{e}^{\text{end}}$, respectively. $N_{p}$ and $N_{e}$ denote the numbers of predicate and entity queries. An example is shown in Fig. 3. Taking $Q_{p}^{\text{end}}\in\mathbb{R}^{N_{e}\times c}$ as an example, the mask vectors $V_{\text{mask}}\in\mathbb{R}^{N_{e}\times 1}$ are generated with a mask ratio $\alpha$. The values of $V_{\text{mask}}$ are set to 0 with probability $\alpha$, and the remaining values are set to $(1/(1-\alpha))$. Then, the dimensions of $V_{\text{mask}}$ are expanded and multiplied with $Q_{p}^{\text{end}}$ to produce $\bar{Q}_{e}^{\text{end}}$. Finally, the mean feature of $\bar{Q}_{e}^{\text{end}}$ is used to match the pre-trained features, as described in Eq. 13.
@@
CLIP-based Classifier After feature distillation, the decoder features $Q_{p}^{end}$ and $Q_{e}^{end}$ are aligned with CLIP. To effectively utilize these features, we introduce classifiers based on CLIP text features. Specifically, we generate descriptions by replacing * in templates like ”A photo of a/an *” or ”A photo of X * Y” with the corresponding entity and predicate classes from the target dataset. These descriptions are input to the CLIP text encoder to obtain the corresponding feature vectors $W_{ent}$ and $W_{pre}$. The resulting vectors are then used to initialize the parameters of the entity and predicate classifiers. Through the classifier layer, the decoder features $Q_{p}^{end}$ and $Q_{e}^{end}$ are used to calculate their cosine similarity with each category, and probabilities are derived from these similarity scores. During training, both classifiers are fine-tuned with a smaller learning rate to improve performance on the SGG datasets. The entire process is illustrated in Fig. 3.
@@ 4.4 Graph Assembling
Inspired by previous work [18], this section elaborates on the process of combining predicted entity and object-aware predicate nodes to generate triplets. Taking the connection between objects and predicates as an example, given the object node set $N_{o}$ and the predicate node set $N_{p}$, we first construct the adjacency matrix $M_{o}$, which represents the distance between these nodes. The calculation is as follows:
$M_{o}=d_{loc}(B_{o},B_{po})\cdot d_{cls}(P_{o},P_{po}),$ (14)
where $d_{loc}$ and $d_{cls}$ are distance functions used to measure matching quality in terms of bounding box locations and classes. Specifically, $B_{o}$, $B_{po}$, $P_{o}$, and $P_{po}$ represent the bounding box predictions and classification distributions for objects and the indicators, respectively. Similarly, the adjacency matrix $M_{s}$ between objects and predicates is calculated in the same way. By selecting the top $K$ relationships from the two adjacency matrices $M_{s}$ and $M_{o}$, the final predictions are obtained in the form $T=\{(B_{o},B_{s},B_{p},P_{o},P_{s},P_{p})\}$. Additional details on graph generation are provided in the supplementary materials.
@@ 4.5 Training and Inference
Training To optimize the parameters of the proposed model, we design a multi-task loss function consisting of three components: $L_{ent}$ for the loss for entity prediction, $L_{pre}$ for predicate prediction, and $L_{mimic}$ for feature distillation. The overall loss function is defined as follows:
@@
Table 1: The Results Comparison in Visual Genome
Method mR@ R@ mR@100
20 50 100 20 50 100 Head Body Tail
Two-stage MOTIFS [35] 4.2 5.7 - 21.4 27.2 - - - -
RelDN [36] - 6.0 7.3 - 31.4 35.9 - - -
VCTree-TDE [29] - 9.3 11.1 - 19.4 23.2 - - -
BGNN [17] 7.5 10.7 12.6 23.3 31.0 35.8 34.0 12.9 6.0
RepSGG [22] 6.7 9.3 11.4 22.5 29.6 34.8 31.3 11.2 5.3
One-stage FCSGG [23] 2.7 3.6 4.2 16.1 21.3 25.1 - - -
SRCNN [30] 6.2 8.6 10.3 26.1 33.5 38.4 - - -
ISGG [12] - 8.0 8.8 - 29.7 32.1 31.7 9.0 1.4
Relationformer [26] 4.6 9.3 10.7 22.2 28.4 31.3 - - -
SGTR [18] - 12.0 15.2 - 24.6 28.4 28.2 18.6 7.1
RelTR [4] 6.8 10.8 12.6 21.2 27.5 - 30.6 14.4 5.0
SG2HOI [9] 8.9 11.4 13.9 21.2 25.9 30.3 27.5 18.2 5.3
DSGG [8] - 13.0 17.3 - 32.9 38.5 - - -
EGTR [10] 5.5 7.9 10.1 23.5 30.2 34.3 - - -
PGSG [19] - 10.5 12.7 - 20.3 23.6 - - -
BCTR (ours) 8.1 13.7 18.4 20.1 24.8 27.7 27.7 22.0 11.9
One-stage with ISGG+Rw [12] - 15.7 17.8 - 27.2 29.8 28.5 18.8 13.3
statistical-based SGTR+Bilvl [18] - 15.8 20.1 - 20.6 25.0 21.7 21.6 17.1
long-tail RelTR+LA [4] 9.7 14.2 - 19.8 25.9 - 28.3 19.4 10.2
strategy BCTR (ours)+LA 12.7 17.4 20.9 17.2 21.9 25.2 24.6 23.4 17.4
Since the entity detector follows a DETR-like architecture, $\mathcal{L}_{ent}$ takes a similar form as described in [1]. To calculate $\mathcal{L}_{pre}$, we first convert the ground-truth relationships in the image into the same format as predictions $T$, denoted as $T_{gt}$. Then, the Hungarian matching algorithm is used to measure the cost between the ground truth and predictions, incorporating both predicate and entity information. The cost is computed as follows:
@@
Inference During inference, the feature distillation module is removed, as it is only used during training. After constructing the graph with predictions, we obtain $K\cdot N_{p}$ triplets. To produce the final results, we first filter out self-connected predictions where the object and subject in a triplet are identical. Next, we re-rank all triplets based on a belief score $S$ and select the top $K$ as the final prediction results. The belief score $S$ is calculated as the product of the classification probabilities of the corresponding subject entity, object entity, and predicate.
@@ 5 Experiments
Datasets We conduct evaluation experiments on two representative scene graph generation datasets, namely Visual Genome [15] and Open Image V6 [16].
@@
Evaluation Metrics The evaluation metrics follow previous works [28]. For Visual Genome, mean Recall (mR@) represents the average recall across all classes, while Recall (R@) reflects the model’s overall recall performance across categories. Additionally, we report mR@100 for each category group (head, body, and tail) to assess the model’s performance on long-tail distributions. For Open Images V6, in addition to R@ and mR@, we use weighted metrics $wmAP_{phr}$, $wmAP_{rel}$ and $score_{wtd}$) to provide a more class-balanced evaluation.
Implementation Details We use ResNet-101 and DETR as the image feature extraction modules. The entity decoder adopts the same architecture as the DETR decoder. The numbers of entity and predicate queries are set to 100 and 160, respectively, and the number of iterations (stages) is empirically set to 6. To ensure training convergence, we begin by training the entity decoder on the dataset with the entity feature distillation loss as defined in Eq. 13. Next, we train the entire model with feature distillation on the dataset, keeping the entity decoder frozen. Finally, we unfreeze the entity decoder and activate bidirectional attention for joint training. Feature distillation losses are calculated for the entity and decoder after BCG. Furthermore, to ensure that BCG’s feature interaction performs in a semantically aligned space, we also compute the entity distillation loss before BCG to enforce constraints on the feature space during joint learning. Additional details on parameter settings are provided in the supplementary materials.
@@ 5.1 Comparisons with State-of-the-Art Methods
Table 2: The Results Comparison in Open Image V6
Method mR@ R@ wmAP wmAP score
50 50 rel phr
HOTR [13] 40.1 52.7 19.4 21.5 26.9
AS-Net [3] 35.2 55.3 25.9 27.5 32.4
SGTR(Li et al. 2022) 42.6 59.9 37.0 38.7 42.3
PGTR [19] 40.7 62.0 22.7 28.0 30.8
BCTR (ours) 48.8 68.6 36.0 39.0 42.5
Table 3: Ablation Study on Model Components
BCG RFA mR@20/50/100 R@20/50/100 H/B/T
7.4/12.8/17.0 19.6/24.2/27.3 26.9/21.3/9.7
7.7/13.2/17.6 19.9/24.6/27.6 27.2/21.5/10.9
7.9/13.2/17.9 20.4/25.1/28.2 28.1/22.4/10.5
8.1/13.7/18.4 20.1/24.8/27.7 27.7/22.0/11.9
@@
Baselines Since BCTR is a one-stage scene graph generation method, we primarily compare it with other current one-stage SGG methods, which are mainly transformer-based [18, 12, 4, 26, 9]. Additionally, we compare our method with several representative two-stage methods [17, 29, 35, 36]. Although existing one-stage methods do not specifically tackle the long-tail problem, previous studies have reported results using statistical-based unbiased training (e.g., bi-level sampling [17]) or inference (e.g., logit adjustment [24]) techniques. For further comparison, we also report our method’s performance with these statistical-based unbiased approaches.

### v3

@@ 4.3 Random Feature Alignment
4.3.1 Random Feature Alignment
As illustrated in Fig. 1 (a), the class imbalance in SGG datasets makes it challenging for BCG to capture feature interaction patterns for tail relationships from rare samples during training. Additionally, the learned feature interactions perform poorly on unseen relationships during inference. To address these issues, we propose RFA, which aligns the SGG feature space with that of a pre-trained vision-language model. In this semantically aligned feature space, BCG can effectively learn feature interaction patterns during training, and the learned patterns generalize better to unseen relationships.
@@ 4.3.1 Random Feature Alignment
Specifically, the input image is fed into the CLIP visual encoder to obtain $\mathbf{V}_{clip}$. Since CLIP is trained with image captions, the encoder features $\mathbf{V}_{clip}$ are inclined to capture the gist of the image. However, images often contain multiple relationships. Simply aligning the decoder’s features $\mathbf{Q}_{p}^{\text{end}}$ and $\mathbf{Q}_{e}^{\text{end}}$ with the CLIP feature $\mathbf{V}_{clip}$ may compromise the feature diversity of SGG. To preserve this diversity, we randomly select a subset of the decoder’s features before each alignment, ensuring that the distilled features retain diversity while being aligned with CLIP. The overall masked distillation process is as follows:
$\mathcal{L}_{mimic}=\left|\mathbf{V}_{clip}-\frac{1}{N_{p}}\sum_{i=1}^{N_{p}}\bar{\mathbf{Q}}_{p}^{\text{end}}\right|+\left|\mathbf{V}_{clip}-\frac{1}{N_{e}}\sum_{i=1}^{N_{e}}\bar{\mathbf{Q}}_{e}^{\text{end}}\right|,$ (13)
where $\bar{\mathbf{Q}}_{p}^{\text{end}}$ and $\bar{\mathbf{Q}}_{e}^{\text{end}}$ are sampled from $\mathbf{Q}_{p}^{\text{end}}$ and $\mathbf{Q}_{e}^{\text{end}}$ via randomly pooling. $N_{p}$ and $N_{e}$ denote the numbers of predicate and entity queries. An example is shown in Fig. 3. Taking $\mathbf{Q}_{p}^{\text{end}}\in\mathbb{R}^{N_{e}\times c}$ as an example, the mask vectors $\mathbf{V}_{mask}\in\mathbb{R}^{N_{e}\times 1}$ are generated with a mask ratio $\alpha$. The values of $\mathbf{V}_{mask}$ are set to 0 with probability $\alpha$, and the remaining values are set to $(1/(1-\alpha))$. Then, the dimensions of $\mathbf{V}_{mask}$ are expanded and multiplied with $\mathbf{Q}_{p}^{\text{end}}$ to produce $\bar{\mathbf{Q}}_{e}^{\text{end}}$. Finally, the mean feature of $\bar{\mathbf{Q}}_{e}^{\text{end}}$ is used to match the pre-trained features, as described in Eq. 13.
@@
4.3.2 CLIP-based Classifier
To effectively utilize the aligned features $\mathbf{Q}_{p}^{\text{end}}$ and $\mathbf{Q}_{e}^{\text{end}}$, we introduce classifiers based on CLIP text features. Specifically, we generate descriptions by replacing * in templates like "A photo of a/an *" or "A photo of X * Y" with all the entity and predicate classes from the dataset. These descriptions are input to the CLIP text encoder to obtain the feature matrices $\mathbf{W}_{\text{ent}}\in\mathbb{R}^{d\times D_{e}}$ and $\mathbf{W}_{\text{pre}}\in\mathbb{R}^{d\times D_{p}}$, where $d$, $D_{e}$, and $D_{p}$ denote the dimension of CLIP features, the number of entity classes, and the number of predicate classes in the dataset, respectively. These matrices are used to initialize the last layer of the entity and predicate classifiers. Through the feature mapping and classifier layers, the decoder features $\mathbf{Q}_{p}^{\text{end}}$ and $\mathbf{Q}_{e}^{\text{end}}$ are projected and then compared with each category embedding via cosine similarity, from which the classification probabilities are computed. During training, both classifiers are fine-tuned with a small learning rate to improve performance on the SGG datasets. The entire process is illustrated in Fig. 3.
@@ 4.4 Graph Assembling
This section outlines the process of generating final triplets using entity nodes $\mathcal{V}_{e}$ and predicate nodes $\mathcal{V}_{p}$. The process begins with the construction of adjacency matrices between $\mathcal{V}_{e}$ and $\mathcal{V}_{p}$, where the matrix elements are determined based on indicators of predicate and entity nodes. Specifically, two adjacency matrices $\mathbf{M}_{s}$ and $\mathbf{M}_{o}$ are computed. $\mathbf{M}_{o}$ represents the relationships between subject nodes and predicate nodes, and $\mathbf{M}_{s}$ represents the relationships between object nodes and predicate nodes. The relationship distances are derived using the corresponding predicted classes and bounding boxes. Taking the pair of objects and predicates as an example, given the object node set $\mathcal{V}_{o}$ and the predicate node set $\mathcal{V}_{p}$, we first construct the adjacency matrix $\mathbf{M}_{o}$, which represents the distance between these nodes. The calculation is as follows:
$\mathbf{M}_{o}=d_{loc}(\mathbf{B}_{o},\mathbf{B}_{po})\cdot d_{cls}(\mathbf{P}_{o},\mathbf{P}_{po}),$ (14)
where $d_{loc}$ and $d_{cls}$ are distance functions used to measure matching quality of bounding box locations and classes. Specifically, $\mathbf{B}_{o}$, $\mathbf{B}_{po}$, $\mathbf{P}_{o}$, and $\mathbf{P}_{po}$ represent the bounding box predictions and classification distributions for objects and the indicators, respectively. Similarly, the adjacency matrix $\mathbf{M}_{s}$ between subjects and predicates is calculated in the same way. For each predicate node in $\mathcal{V}_{p}$, we select the top $k$ edges from the two adjacency matrices $\mathbf{M}_{s}$ and $\mathbf{M}_{o}$. Consequently, a total of $N_{p}\times k$ candidate triplets are generated as predictions in the form of $T=\{(\mathbf{B}_{o},\mathbf{B}_{s},\mathbf{B}_{p},\mathbf{P}_{o},\mathbf{P}_{s},\mathbf{P}_{p})\}$, where $\mathbf{B}_{o},\mathbf{B}_{s},\mathbf{B}_{p}$ and $\mathbf{P}_{o},\mathbf{P}_{s},\mathbf{P}_{p}$ represent the bounding boxes and class predictions of its object, subject, and predicate, respectively.
@@ 4.5 Training and Inference
4.5.1 Training
To optimize the parameters of the proposed model, we design a multi-task loss function consisting of three components: $\mathcal{L}_{ent}$ for the entity prediction, $\mathcal{L}_{pre}$ for predicate prediction, and $\mathcal{L}_{mimic}$ for feature distillation. The overall loss function is defined as follows:
@@ 4.5.1 Training
Since the entity detector follows a DETR-like architecture, $\mathcal{L}_{ent}$ takes a similar form as described in [1]. To calculate $\mathcal{L}_{pre}$, we first convert the ground-truth relationships in the image into the same format as predictions $T$, denoted as $T_{gt}$. Then, the Hungarian matching algorithm measures the cost between the ground truth and predictions, incorporating both predicate and entity information. The cost is computed as follows:
@@
4.5.2 Inference
During inference, the feature distillation module is removed, as it is used only during training. After constructing the graph with predictions, we obtain $k\cdot N_{p}$ triplets. To produce the final results, we first filter out self-connected predictions where the object and subject in a triplet are identical. Next, we re-rank all triplets based on a belief score $S$ and select the top $k$ as final predictions. The belief score $S$ is calculated as the product of the classification probabilities of the corresponding subject entity, object entity, and predicate.
@@ 5 Experiments
5.1 Datasets and Evaluation Metrics
5.1.1 Visual Genome
@@
Visual Genome [17] is the most representative SGG dataset, consisting of 57k training images, 5k validation images, and 26k test images. The dataset includes 150 object categories and 50 predicate categories. Following previous one-stage SGG studies [34, 11], we adopt the Scene Graph Detection (SGDET) setting to evaluate the proposed model. Following Tang et al. [37], we use mean Recall@k (mR@k) and Recall@k (R@k) as the evaluation metrics. mR@k represents the average recall across all classes, while R@k reflects the model’s overall recall performance across categories. Inspired by previous work [46, 31], we use the overall metric F@K to jointly evaluate R@K and mR@K, which is the harmonic average of R@K and mR@K. Additionally, we report mR@100 for each category group (head, body, and tail) to assess the model’s performance on long-tail distributions.
5.1.2 Open Images V6
We also conduct experiments on the larger-scale Open Image V6 [19] dataset, which consists of 126k training images, 5.3k test images, and 1.8k validation images. The dataset includes 288 object categories and 30 predicate categories. In addition to R@k, mR@k and F@k, we also report weighted mean average precision (AP) of phrase detection $wmAP_{phr}$ and relationship detection $wmAP_{rel}$. The $score_{wtd}$ is computed from R@k, $wmAP_{phr}$ and $wmAP_{rel}$ for a more class-balanced evaluation [4].
5.2 Implementation Details
We use ResNet-101 and DETR as the image feature extraction modules. The entity decoder adopts the same architecture as the DETR decoder. For Visual Genome [17], the numbers of entity nodes and predicate nodes are 100 and 160, respectively. For OpenImage V6 [19], these numbers are 100 and 180. The number of iterations (stages) is empirically set to 6. The value of $k$ is set to 25 for Visual Genome and 40 for OpenImage V6.
To ensure stable training and faster convergence, we first train the model using RFA. The BCG is then incorporated into the training process for joint optimization. To train the model with RFA, we employ the AdamW optimizer with a weight decay rate of $1\times 10^{-4}$. The entire model includes two fully connected layers followed by entity and predicate decoders, which map features into object and predicate category dimensions. The training process begins by optimizing the entity decoder on the target dataset using the entity feature distillation loss. Subsequently, the model is trained jointly with the feature distillation loss. The parameters of the entity decoder are frozen, while the predicate decoder remains trainable with a learning rate of $1\times 10^{-4}$. Additionally, classifier parameters initialized from CLIP text features are updated with a small learning rate of $1\times 10^{-5}$. The training process uses a batch size of 24 and runs for 90 epochs on 4 NVIDIA V100 GPUs. The loss function incorporates coefficients of 20, 1.0, and 0.5 for the RFA loss, entity detection loss, and predicate detection loss, respectively.
During joint training, the BCG is integrated into the model. Specifically, an additional entity branch and a BiAtt module are introduced to enhance feature interaction. The entity decoder in BCG shares parameters with the entity decoder used during feature extraction, and these shared parameters are updated jointly during optimization. To ensure that BCG’s feature interaction occurs in a semantically aligned space, we also compute the entity distillation loss at the entity decoder, enforcing constraints on the feature space during joint learning. The optimizer, learning rate, and batch size remain the same as in the first stage. The model is trained for 90 epochs to achieve final optimization.
Table 1: The Results Comparison in Visual Genome.
Method mR@ R@ mR@100 F@
20 50 100 20 50 100 Head Body Tail 20 50 100
Two-stage MOTIFS [44] 4.2 5.7 - 21.4 27.2 - - - - 7.0 9.4 -
RelDN [48] - 6.0 7.3 - 31.4 35.9 - - - - 10.1 12.1
VCTree-TDE [36] - 9.3 11.1 - 19.4 23.2 - - - - 12.6 15.0
BGNN [22] 7.5 10.7 12.6 23.3 31.0 35.8 34.0 12.9 6.0 11.3 15.9 18.6
RepSGG [27] 6.7 9.3 11.4 22.5 29.6 34.8 31.3 11.2 5.3 10.3 14.2 17.2
One-stage FCSGG [28] 2.7 3.6 4.2 16.1 21.3 25.1 - - - 4.6 6.2 7.2
SRCNN [38] 6.2 8.6 10.3 26.1 33.5 38.4 - - - 10.0 13.7 16.2
ISGG [14] - 8.0 8.8 - 29.7 32.1 31.7 9.0 1.4 - 12.6 13.8
Relationformer [34] 4.6 9.3 10.7 22.2 28.4 31.3 - - - 7.6 14.0 15.9
SGTR [20] - 12.0 15.2 - 24.6 28.4 28.2 18.6 7.1 - 16.1 19.8
RelTR [4] 6.8 10.8 12.6 21.2 27.5 - 30.6 14.4 5.0 10.3 15.5 -
EGTR [11] 5.5 7.9 10.1 23.5 30.2 34.3 - - - 8.9 12.5 15.6
PGSG [21] - 10.5 12.7 - 20.3 23.6 - - - - 13.8 16.5
BCTR (ours) 8.1 13.7 18.4 20.1 24.8 27.7 27.7 22.0 11.9 11.5 17.6 22.1
One-stage with ISGG+Rw [14] - 15.7 17.8 - 27.2 29.8 28.5 18.8 13.3 - 19.9 22.3
statistical-based SGTR+Bilvl [20] - 15.8 20.1 - 20.6 25.0 21.7 21.6 17.1 1.0 17.9 22.3
long-tail RelTR+LA [4] 9.7 14.2 - 19.8 25.9 - 28.3 19.4 10.2 13.0 18.3 -
strategy BCTR (ours)+LA 12.7 17.4 20.9 17.2 21.9 25.2 24.6 23.4 17.4 14.6 19.4 22.8
[…]
Method R@50 $wmAP_{rel}$ $wmAP_{phr}$ $score_{wtd}$ mR@50 F@50
RelDN [48] 72.8 29.9 30.4 38.7 36.8 48.9
GPS-Net [26] 74.7 32.8 33.9 41.6 38.9 51.2
BGNN [22] 74.9 31.2 31.4 40.0 39.4 51.6
HOTR [15] 52.7 19.4 21.5 26.9 40.1 45.5
AS-Net [3] 55.3 25.9 27.5 32.4 35.2 43.0
SGTR [20] 71.0 35.5 38.8 42.4 44.1 54.4
PGSG [21] 62.0 19.7 27.8 28.7 40.7 49.1
BCTR (ours) 68.6 36.0 39.0 43.7 48.8 57.0
@@ 5.3 Comparisons with State-of-the-Art Methods
5.3.1 Baselines
@@
Since BCTR is a one-stage SGG method, we primarily compare it with other current one-stage methods, which are mainly transformer-based [20, 14, 4, 34, 10]. Additionally, we compare our method with several representative two-stage methods [22, 36, 44, 48]. Although existing one-stage methods do not specifically tackle the long-tail problem, previous studies have reported results using statistical-based unbiased training (e.g., bi-level sampling [22]) or inference (e.g., logit adjustment [30]) techniques. For further comparison, we also report our method’s performance with these statistical-based unbiased approaches.
5.3.2 Results of Visual Genome